A Company is upgrading office technology by purchasing inkjet printers. LCD monitors, and additional memory chips. The total number of pieces of hardware purchased is 48. The cost of each inkjet printer is $119, the cost of each LCD monitor is $138, and the cost of each memory chip is $93. The total amount of money spent on new hardware came to $5283. They purchased two times as many memory chips as they did LCD monitors. Determine the number of each that was purchased.
linear system of equations is
119x + 138y + 93z = 5283
x + y + z = 48
-2y + z = 0
find the number of each that was purchased.

Answer by ikleyn(52848)

.
A Company is upgrading office technology by purchasing inkjet printers. LCD monitors, and additional memory chips.
The total number of pieces of hardware purchased is 48. The cost of each inkjet printer is $119, the cost of each LCD monitor is $138,
and the cost of each memory chip is $93. The total amount of money spent on new hardware came to $5283.
They purchased two times as many memory chips as they did LCD monitors. Determine the number of each that was purchased.
~~~~~~~~~~~~~~~~~~~~~

Let M be the number of monitors.
Then the number of memory chips is 2M.

Let P be the number of printers.

Equations:

M + 2M + P = 48,
119P + 138M + 93*(2M) = 5283.

Or

  3M +    P =   48,   (1)
324M + 119P = 5283.   (2)

Solve on your own by any method you know.